OpenStudy (anonymous):

6x^2+13x-20=0 solve this as a quadratic equation

OpenStudy (anonymous):

\[-b \pm \sqrt{b^2 - 4ac} \over 2a\] \[-13 \pm \sqrt{13^2 - 4*6*-20} \over 2*6\] x = \[-13 \pm \sqrt{649} \over 12\]

OpenStudy (mimi_x3):

use the quadratic formula x = 1.0396 , x = -2.39

OpenStudy (anonymous):

nice!!!! thaks for the answer!!! ^_^